# ToMoE: Converting Dense Large Language Models to Mixture-of-Experts through Dynamic Structural Pruning

### v1

@@ 3.3 MHA top-K Routing
LLM-Pruner (Ma et al., 2023) 13.56 17.90 31.05 12.19 19.56 32.20
LLM Surgeon (van der Ouderaa et al., 2024) 7.83 10.39 15.38 6.21 7.25 9.43
ShortGPT (Men et al., 2024) 33.21 71.04 268.11 30.48 48.83 187.23
SLEB (Song et al., 2024) 11.23 29.10 103.38 8.24 11.76 27.67
SliceGPT (Ashkboos et al., 2024) 10.47 15.19 24.82 8.68 12.56 20.57
ModeGPT (Lin et al., 2024) 7.51 8.41 11.88 6.10 6.95 8.95
DISP-LLM (Gao et al., 2024) 6.85 8.11 9.84 5.77 6.59 7.11
@@
3.4 Regularizations for MoE Constructions
In Sec. 3.2 and Sec. 3.3, we briefly introduced the design space for constructing MoE models using dynamic structural pruning. In this subsection, we will introduce regularizations tailored to the needs of MoE models.
where $\bigcup$ is the union operator, and $\mathbf{u}$ is the union of all kept positions for each token. For MLP layers, it can be calculated similarly. We than push $\frac{\sum\mathbf{u}}{|\mathbf{u}|}$ ($|\mathbf{u}|$ represents the size of $\mathbf{u}$) to 1:
${}\mathcal{R}_{\text{U}}=\frac{1}{L}\sum_{l=1}^{L}f(\frac{\sum\mathbf{u}_{l}}{|\mathbf{u}_{l}|},1),$ (7)
where $f(\cdot,\cdot)$ can be any regression loss functions, and we will choose $f$ later.
Parameter Regularization. For a sparse MoE model, we also need to control the number of active parameters and the number of active parameters should be fixed. To achieve this goal, we can directly penalize the maximum width from different layers. We choose the maximum width over experts instead of the mean, median, or other alternatives because the maximum provides precise control over the upper bound of the maximum number of active parameters.
[…]
${}\mathcal{R}_{\text{P}}=f(\text{T}(\mathbf{d_{\text{MoE}}}),p\text{T}_{\text{total}}),$ (8)
where $\text{T}_{\text{total}}$ is the total number of parameters, and $p\in(0,1]$ represents the ratios of the active parameters. For $f$ in Eq. 8 and Eq. 7, we use the following function $f$:
$f(x,y)=\log(\max(x,y)/\min(x,y))$
Load Balancing Regularization. When we try to find the configurations of experts, we also apply the load balancing regularization to encourage a balanced load across experts (Lepikhin et al., 2021a; Fedus et al., 2022). We follow the load balancing loss from the Switch Transformer (Fedus et al., 2022):
${}\mathcal{R}_{\text{L}}=N\sum_{i=1}^{N}F_{i}P_{i},$ (9)
here, $F_{i}=\frac{1}{T}\sum_{t=1}^{T}\mathbbm{1}(\mathbf{G}_{t,i}=1)$, where $\mathbbm{1}(\cdot)$ is an indicator function and that returns $1$ if the condition is true, and 00 otherwise. $F_{i}$ represents the fraction of tokens assigned to the $i$-th expert. $P_{i}=\frac{1}{T}\sum\sigma_{s}({G}_{t,i})$, where ${G}=\text{Router}(\mathbf{X})$ is the router output before ST-G, and $\sigma_{s}$ is the softmax function again. $P_{i}$ is the fraction of the router probability allocated for the $i$-th expert. $\mathcal{R}_{\text{L}}$ will encourages uniform routing across different experts as shown in (Fedus et al., 2022). The combination of Eq. 8 and Eq. 9 creates an interesting phenomenon where they encourage uniform allocation of width among experts.
@@ 3.4 Regularizations for MoE Constructions
3.5 Learning to Construct MoEs
Based on the aforementioned techniques, we can construct MoEs from the dense LLMs by training router parameters, projection parameters, and hypernetwork parameters. At the same time, we freeze all the parameters in the original model. Doing so allows us to quickly construct an effective MoE model with a resource budget comparable to structural pruning. The overall framework of our method is shown in Fig. 2. The overall training objective function can be formulated as:
where $\Theta=[\Theta_{\text{HN}},\Theta_{\text{Router}},\Theta_{\text{Proj-MHA}},\Theta_{\text{Proj-MLP}}]$, $\Theta_{\text{HN}}$ is trainable parameters for the hypernetwork in Eq. 1, $\Theta_{\text{Router}}$ and $\Theta_{\text{Proj-MLP}}$ are trainable parameters for the Router and the project module in Eq. 3, $\Theta_{\text{Proj-MHA}}$ is the trainable parameters of the projection modules given in Eq. 5, $\mathcal{R}_{\mathbf{P}}$, $\mathcal{R}_{\mathbf{U}}$, and $\mathcal{R}_{\mathbf{L}}$ are regularization terms defined in Sec. 3.4, and $\alpha$, $\beta$ and $\gamma$ are hyperparameters to control the strength of these regularization terms. We use $f$ to represent the original dense model and use $f^{\prime}$ as the model equipped with our designed modules for MoE construction. Under this setting, we use $\mathcal{L}(\cdot,\cdot)$ to calculate the KL divergence between the logits of $f$ and $f^{\prime}$ and use it as the guidance to preserve the capacity of the dense model (Hinton et al., 2015). We also found that only using the KL divergence leads to the best performance, and this observation complies with the experimental setup in (Muralidharan et al., 2024). Also, note that we perform in-place knowledge distillation since the original model weights are frozen. Thus the knowledge distillation process does not introduce overheads in terms of GPU memory.
After learning how to construct the MoE, we prune the MLP layer to $N$ experts with shared weights. After pruning the MLP layer, we save $\frac{1}{N}\mathbf{1}^{\top}\mathbf{E}$ for MHA layers as the bias of the $\text{Proj}_{\text{E}}^{\text{\tiny MHA}}$, and we drop $\text{Proj}_{\text{D}}^{\text{\tiny MLP}}$ and convert Eq. 5 into a Top-k routing function as well as use $s_{0}$ for pruning $\mathbf{W}_{Q}$ and $\mathbf{W}_{K}$. Our construction also enables converting the MoE back into a pseudo-MoE model. The MoE model and the pseudo-MoE model are equivalent, and more details are given in the Appendix B.3.
@@ 3.5 Learning to Construct MoEs
70% ShortGPT (Men et al., 2024) 48.65 32.85 64.31 64.33 56.13 53.25
SliceGPT (Ashkboos et al., 2024) 58.88 33.36 68.55 58.01 49.86 53.73
LLM Surgeon (van der Ouderaa et al., 2024) 63.09 36.69 73.56 61.09 60.72 59.03
DISP-LLM  (Gao et al., 2024) 59.81 33.19 71.82 62.27 63.43 58.10
DISP-LLM Alpaca (Gao et al., 2024) 60.10 37.03 73.72 63.93 62.87 59.53
ModeGPT (Lin et al., 2024) 63.26 38.73 70.40 67.32 63.26 60.78
ModeGPT-Alpaca (Lin et al., 2024) 65.49 39.16 73.34 66.22 65.90 62.02
60% ShortGPT (Men et al., 2024) 41.16 29.94 60.12 60.46 43.67 47.07
SliceGPT (Ashkboos et al., 2024) 36.49 24.57 54.90 53.43 34.80 40.84
LLM Surgeon (van der Ouderaa et al., 2024) 52.31 30.29 69.26 54.38 48.04 50.86
ModeGPT (Lin et al., 2024) 49.45 30.03 64.96 61.96 53.01 51.88
ModeGPT-Alpaca (Lin et al., 2024) 59.76 34.73 70.35 64.40 58.63 57.58
@@
50% LLM Surgeon  (van der Ouderaa et al., 2024) 44.91 26.28 64.36 52.57 40.29 45.68
DISP-LLM  (Gao et al., 2024) 43.06 25.85 63.93 54.54 63.43 46.72
DISP-LLM Alpaca (Gao et al., 2024) 51.14 30.20 68.34 56.20 49.35 51.05
ToMoE (Ours) 56.65 33.87 71.00 58.56 60.26 56.07
@@
75% ShortGPT-Alpaca (Men et al., 2024) 38.13 31.40 60.94 54.22 31.52 43.24
SliceGPT-Alpaca (Ashkboos et al., 2024) 44.44 29.27 57.56 58.48 41.08 46.17
ModeGPT-Alpaca (Lin et al., 2024) 67.05 41.13 75.52 69.61 66.49 63.96
@@ 4.1 Settings
Models. We evaluate our ToMoE method using several LLMs with decoder blocks. Specifically, we choose the following models: LLaMA-2 (Touvron et al., 2023b): LLaMA-2 7B and LLaMA-2 13B; LLaMA-3 8B (Dubey et al., 2024); Phi-2 (Javaheripi et al., 2023) (shown in Appendix D); Qwen-2.5 (Yang et al., 2024): Qwen-2.5 7B and Qwen-2.5 14B.
@@
Implementations. We implemented ToMoE using Pytorch (Paszke et al., 2019) and Hugging Face transformer library (Wolf et al., 2020). We freeze the model weights when training the modules with learnable parameters $\Theta$ in Eq. in Obj. 10. We use the AdamW (Loshchilov & Hutter, 2019) optimizer to optimize $\Theta$. $\Theta$ is trained for 10,000 iterations for all models. For all experiments, we set $\alpha=16$, $\beta=2.0$, and $\gamma=1.0$, where $\alpha$, $\beta$, and $\gamma$ are defined in Obj. 10. Without specific descriptions, the number of experts for ToMoE is $8$ across all settings. Depending on the size of the base model, we use 1 to 4 NVIDIA A100 GPUs to train $\Theta$. More implementation details can be found in the Appendix C.
@@
Datasets. We provide two training settings for all modules with learnable parameters $\Theta$: (1) using WikiText, and (2) using a mixed dataset comprising WikiText, Alpaca, and Code-Alpaca (mixing ratio: 1:1:1). Based on our understanding, ToMoE benefits from a diverse mixture of datasets to effectively construct experts. Following previous methods (Ashkboos et al., 2024; Gao et al., 2024), we evaluate our method and other methods on five well-known zero-shot tasks: PIQA (Bisk et al., 2020); WinoGrande (Sakaguchi et al., 2021); HellaSwag (Zellers et al., 2019); ARC-e and ARC-c (Clark et al., 2018). We use llm-eval-harness (Gao et al., 2021) to evaluate the compressed models.
Baselines. We compare ToMoE across baselines from structural pruning like LLM-Pruner (Ma et al., 2023), SliceGPT (Ashkboos et al., 2024), ShortGPT (Men et al., 2024), SLEB (Song et al., 2024), LLMSurgeon (van der Ouderaa et al., 2024), ModeGPT (Lin et al., 2024), and DISP-LLM (Gao et al., 2024). We also include semi-structure pruning baselines like SparseGPT (Frantar & Alistarh, 2023), Wanda (Sun et al., 2024), and Pruner-Zero (Dong et al., 2024b).
@@ 4.2 Language Modeling
Figure 5: (a) Model width and the union of all experts. (b) Costs of different learning-based methods.
Figure 6: Ablation Study on Qwen-2.5 7B.
@@
Tab. 1 presents the perplexity results of structured pruning methods applied to LLaMA-2 models of sizes 7B and 13B on the WikiText-2 dataset, comparing various methods with 70%, 60%, and 50% of active parameters which correspond to pruning ratios of 30%, 40%, and 50% for pruning. Across all pruning ratios, ToMoE consistently achieves the lowest perplexity compared to other methods, even outperforming many approaches with significantly larger numbers of active parameters. For instance, ToMoE with 50% active parameters achieves a perplexity of 8.36, which is superior to LLM-Pruner, ShortGPT, SLEB, and SliceGPT at a 30% pruning ratio. Furthermore, ToMoE with 50% active parameters surpasses ModeGPT and LLM Surgeon at a 40% pruning ratio. While the gap between ToMoE and DISP-LLM is smaller, it is still obvious at a 50% pruning ratio: ToMoE achieves a perplexity that is 1.48 points lower than DISP-LLM. ToMoE also exhibits superior performance with the LLaMA-2 13B model, maintaining a similar advantage over other methods as observed with the LLaMA-2 7B model. This demonstrates the effectiveness of ToMoE in maintaining strong language modeling performance, even with much fewer active parameters. Tab. 2 presents a comparison of our method against semi-structural pruning techniques. Our approach consistently achieves the lowest perplexity while retaining 50% of the active parameters. Moreover, the performance gap between our method and the semi-structural pruning methods is also obvious. On the LLaMA-2 7B model, SparseGPT achieves the second-best performance, with our method improving upon it by 1.81 in terms of perplexity. For the LLaMA-2 13B model, Pruner-Zero shows the second-best performance, while ToMoE further reduces the perplexity by 0.63. The comparison against seme-structural pruning methods further demonstrates the advantage of our method on the language modeling task.
4.3 Zero-Shot Performance
@@
In Tab. 3, we present the zero-shot performance of various methods on LLaMA-2 7B, LLaMA-3 8B, and Qwen-2.5 7B. Our method consistently achieves the best average performance across all models. Compared to relatively weaker methods like ShortGPT and SliceGPT, our approach demonstrates higher average performance (ToMoE 50%: 56.07 vs. SliceGPT 70%: 53.73 and ShortGPT 70%: 53.25) while utilizing 20% fewer active parameters on the LLaMA-2 7B model. Against stronger baselines such as LLM Surgeon and DISP-LLM, ToMoE achieves better performance with 10% fewer active parameters on the same model. For instance, ToMoE (50%) outperforms LLM Surgeon (60%), and ToMoE (60%) outperforms both LLM Surgeon (70%) and DISP-LLM (70%). Although ModeGPT performs closer to ToMoE, the gap remains significant. With 60% active parameters, ToMoE is 3.14 better than ModeGPT. In addition, when keeping 50% active parameters, ToMoE outperforms DISP-LLM and LLM Surgeon by 5.02 and 10.39, respectively. The performance advantage of ToMoE is even larger on LLaMA-3 8B, where it reduces 5% more active parameters than ModeGPT while still achieving a 2.71 performance gain. Furthermore, when removing 15% more active parameters compared to ShortGPT and SliceGPT, ToMoE exceeds their average performance by 18.93 and 16 points, respectively.
For Qwen-2.5 7B, ToMoE significantly outperforms DISP-LLM, consistent with previous findings on other models. We further investigate the effect of the number of experts $N$ when 40% to 50% of the parameters are active. The results indicate that increasing the number of experts to 16 is beneficial. However, further increasing $N$ to 24 provides only marginal or no improvement, likely because a too-large number of experts burdens the learning process.

### v2

@@ 3.4 Regularizations for MoE Constructions
LLM-Pruner (ma2023llm) 13.56 17.90 31.05 12.19 19.56 32.20
LLM Surgeon (van2023llm) 7.83 10.39 15.38 6.21 7.25 9.43
ShortGPT (men2024shortgpt) 33.21 71.04 268.11 30.48 48.83 187.23
SLEB (songsleb) 11.23 29.10 103.38 8.24 11.76 27.67
SliceGPT (ashkboos2023slicegpt) 10.47 15.19 24.82 8.68 12.56 20.57
ModeGPT (lin2024modegpt) 7.51 8.41 11.88 6.10 6.95 8.95
DISP-LLM (gaodisp) 6.85 8.11 9.84 5.77 6.59 7.11
@@
3.5 Learning to Construct MoEs
Based on the aforementioned techniques, MoEs can be constructed from the dense LLMs by training router parameters, projection parameters, and hypernetwork parameters, while keeping all original model parameters frozen. This approach enables the rapid construction of an effective MoE model with a resource budget comparable to that of structural pruning. The overall framework of our method is shown in Fig. 2, and the corresponding training objective function can be formulated as:
@@ 3.5 Learning to Construct MoEs
where $\Theta=[\Theta_{\text{HN}},\Theta_{\text{Router}},\Theta_{\text{Proj-MHA}},\Theta_{\text{Proj-MLP}}]$, $\Theta_{\text{HN}}$ is trainable parameters for the hypernetwork in Eq. 1, $\Theta_{\text{Router}}$ and $\Theta_{\text{Proj-MLP}}$ are trainable parameters for the router and the project module in Eq. 3, $\Theta_{\text{Proj-MHA}}$ is the trainable parameters of the projection modules given in Eq. 5, $\mathcal{R}_{\mathbf{P}}$, $\mathcal{R}_{\mathbf{U}}$, and $\mathcal{R}_{\mathbf{L}}$ are regularization terms defined in Sec. 3.4. And $\alpha$, $\beta$, and $\gamma$ are hyperparameters to control the strength of these regularization terms. Here,$f$ represents the original dense model, and $f^{\prime}$ is the model equipped with our designed modules for MoE construction. Under this setting, we use $\mathcal{L}(\cdot,\cdot)$ to calculate the KL divergence between the logits of $f$ and $f^{\prime}$, which is used as the guidance to preserve the capacity of the dense model (hinton2015distillingknowledgeneuralnetwork). We also found that using the KL divergence alone can lead to the best performance, and this observation complies with the experimental setup in (muralidharan2024compact). Also, note that we perform in-place knowledge distillation since the original model weights are frozen. Thus, the knowledge distillation process does not introduce overheads in terms of GPU memory.
After learning how to construct the MoE, we convert the MLP layer to $N$ experts with shared weights. After pruning the MLP layer, we save $\frac{1}{N}\mathbf{1}^{\top}\mathbf{E}$ for MHA layers as the bias of the $\text{Proj}_{\text{E}}^{\text{\tiny MHA}}$, and we drop $\text{Proj}_{\text{D}}^{\text{\tiny MLP}}$ and convert Eq. 5 into a Top-K routing function as well as use $\mathbf{s}_{0}$ for pruning $\mathbf{W}_{Q}$ and $\mathbf{W}_{K}$. Our construction also enables converting the MoE back into a pseudo-MoE model. The MoE model and the pseudo-MoE model are equivalent, and more details can be found in Appendix B.3.
@@
60% ShortGPT (men2024shortgpt) 41.16 29.94 60.12 60.46 43.67 47.07
SliceGPT (ashkboos2023slicegpt) 36.49 24.57 54.90 53.43 34.80 40.84
LLM Surgeon (van2023llm) 52.31 30.29 69.26 54.38 48.04 50.86
ModeGPT (lin2024modegpt) 49.45 30.03 64.96 61.96 53.01 51.88
ModeGPT-Alpaca (lin2024modegpt) 59.76 34.73 70.35 64.40 58.63 57.58
@@
75% ShortGPT-Alpaca (men2024shortgpt) 38.13 31.40 60.94 54.22 31.52 43.24
SliceGPT-Alpaca (ashkboos2023slicegpt) 44.44 29.27 57.56 58.48 41.08 46.17
ModeGPT-Alpaca (lin2024modegpt) 67.05 41.13 75.52 69.61 66.49 63.96
@@ 4.1 Settings
Models. Our ToMoE method is evaluated using several LLMs with decoder blocks. Specifically, we choose the following models: LLaMA-2 (touvron2023llama2): LLaMA-2 7B and LLaMA-2 13B; LLaMA-3 8B (dubey2024llama); Phi-2 (javaheripi2023phi); Qwen-2.5 (yang2024qwen2): Qwen-2.5 7B and Qwen-2.5 14B. Results for LLaMA-2 13B and Qwen-2.5 14B are presented in the Appendix D.
@@
Implementations. ToMoE is implemented by Pytorch (paszke2019pytorch) and Hugging Face transformer library (wolf-etal-2020-transformers). The model weights are frozen when training the modules with learnable parameters $\Theta$ in Obj. 10. We use the AdamW (loshchilov2018decoupled) optimizer to optimize $\Theta$, which is trained for 10,000 iterations for all models. For all experiments, we set $\alpha=16$, $\beta=2.0$, and $\gamma=1.0$, where $\alpha$, $\beta$, and $\gamma$ are defined in Obj. 10. Without specific descriptions, the number of experts for ToMoE is $8$ across all settings. Depending on the size of the base model, 1 to 4 NVIDIA A100 GPUs are used to train $\Theta$. More implementation details can be found in Appendix C.
@@
Datasets. Two training settings are provided for all modules with learnable parameters $\Theta$: (1) using WikiText merity2016pointer, and (2) using a mixed dataset comprising WikiText, Alpaca alpaca, and Code-Alpaca codealpaca (mixing ratio: 1:1:1). Based on our observation, ToMoE benefits from a diverse mixture of datasets to effectively construct experts. Following previous methods (ashkboos2023slicegpt; gaodisp), our method and other methods are evaluated on five well-known zero-shot tasks: PIQA (bisk2020piqa); WinoGrande (sakaguchi2021winogrande); HellaSwag (zellers2019hellaswag); ARC-e and ARC-c (clark2018think). We further evaluate our method on the following tasks and configurations to ensure consistency with comparison baselines: 32-shot BoolQ (clark2019boolq), SciQ (SciQ), 5-shot WinoGrande, 25-shot ARC-c, 10-shot HellaSwag, TruthfulQA lin2022truthfulqa, and 5-shot MMLU (hendryckstest2021). We use llm-eval-harness (eval-harness) to evaluate the compressed models.
Baselines. ToMoE is compared to baselines from structural pruning methods (ma2023llm; ashkboos2023slicegpt; men2024shortgpt; songsleb; van2023llm; lin2024modegpt; gaodisp), semi-structural pruning methods (frantar2023sparsegpt; sun2024a; dongpruner) and MoE construction methods (zhu2024llama; lee2024breaking; qu2024llama; pei2025cmoe).
Table 4: Compassion against MoE construction methods.
acc-norm acc-norm acc-norm acc acc-norm
LLaMA-2 7B 100% Dense 74.58 46.25 79.11 69.06 75.99 69.00
50% LLaMA-MoE E16A4 (zhu2024llama) 27.02 29.53 49.13 49.49 26.19 36.27
+ fine-tuning 35.52 26.88 49.13 55.71 34.07 40.55
LLaMA-MoE E8A2 (zhu2024llama) 25.76 27.22 50.92 47.51 26.02 35.49
+ fine-tuning 37.71 27.56 58.11 51.78 36.37 42.31
ToMoE (Ours) 56.65 33.87 71.00 58.56 60.26 56.07
Phi-2 100% Dense 78.24 54.01 79.11 75.61 73.86 72.17
78% G-MoEfication (lee2024breaking) 66.46 38.91 70.29 65.75 57.90 59.86
70% G-MoEfication (lee2024breaking) 57.95 35.24 64.80 60.30 49.07 53.47
ToMoE (Ours) 70.79 43.86 77.00 66.38 62.68 64.16
Table 5: Comparison against MoE construction methods on LLaMA-2 7B.
Active Parameters Method # Tokens BoolQ (32) SciQ PIQA WinoG. (5) ARC-C (25) HellaS. (10) Average
acc acc acc acc acc-norm acc-norm
100% LLaMA-2 7B 2T 82.04 90.80 78.78 73.95 53.15 78.55 76.21
50% LLaMA-MoE E8A2 (zhu2024llama) 1.2B 37.83 20.00 49.73 50.12 25.79 26.18 34.27
LLaMA-MoE-v2 E8A2 (qu2024llama) 1.2B 51.25 67.00 56.64 52.88 25.68 35.10 48.59
CMoE E8S1A1 (pei2025cmoe) 0 46.09 65.30 52.77 48.70 23.80 30.12 44.13
+ fine-tuning 1.2B 55.04 77.50 57.12 54.06 27.56 38.79 51.68
ToMoE (Ours) 0.02B 62.42 81.70 71.00 60.41 34.39 60.54 61.39
Table 6: Comparison against LLaMA-MoE-v2 on LLaMA-3 8B.
Active Parameters Method # Tokens BoolQ (32) SciQ PIQA ARC-C (25) TruthfulQA HellaS. (10) MMLU (5) Average
acc acc-norm acc acc-norm acc acc-norm acc
100% LLaMA-3 8B 15T 83.00 93.20 78.51 61.86 51.71 78.79 67.22 73.33
50% LLaMA-MoE-v2 E8A2 (qu2024llama) 7B 74.62 90.60 69.26 42.83 45.62 58.95 37.41 59.61
LLaMA-MoE-v2 E8A1S1 (qu2024llama) 7B 76.88 88.80 67.90 40.19 46.85 53.67 40.89 59.31
ToMoE (Ours) 0.02B 63.67 87.50 72.63 37.29 40.48 63.91 36.31 57.40
@@ 4.2 Language Modeling
Figure 4: Experts token allocation of ToMoE for the LLaMA-3 8B model collected on the WikiText dataset.
@@
Figure 6: (a) Model width and union of experts. (b) Costs of different learning-based methods. (c) Ablation study on Qwen-2.5 7B.
@@
Tab. 1 presents the perplexity results of structured pruning methods applied to LLaMA-2 models of sizes 7B and 13B on the WikiText-2 dataset, comparing various methods with 70%, 60%, and 50% of active parameters setting—corresponding to pruning ratios of 30%, 40%, and 50% for pruning, respectively. Across all pruning ratios, ToMoE consistently achieves the lowest perplexity compared to other methods, even outperforming many approaches with significantly larger numbers of active parameters. For instance, ToMoE with 50% active parameters achieves a perplexity of 8.36, which is superior to LLM-Pruner, ShortGPT, SLEB, and SliceGPT at a 30% pruning ratio. Furthermore, ToMoE with 50% active parameters surpasses ModeGPT and LLM Surgeon at a 40% pruning ratio. While the gap between ToMoE and DISP-LLM is smaller, it is still obvious at a 50% pruning ratio: ToMoE achieves a perplexity that is 1.48 points lower than DISP-LLM. ToMoE also exhibits superior performance with the LLaMA-2 13B model, maintaining a similar advantage over other methods as observed with the LLaMA-2 7B model. This demonstrates the effectiveness of ToMoE in maintaining strong language modeling performance, even with much fewer active parameters. Tab. 2 presents a comparison of our method against semi-structural pruning techniques. Our approach consistently achieves the lowest perplexity while retaining 50% of the active parameters. Moreover, the performance gap between our method and the semi-structural pruning methods is also obvious. On the LLaMA-2 7B model, SparseGPT achieves the second-best performance, with our method improving upon it by 1.81 in terms of perplexity. For the LLaMA-2 13B model, Pruner-Zero shows the second-best performance, while ToMoE further reduces the perplexity by 0.63. The comparison against semi-structural pruning methods further demonstrates the advantage of our method on the language modeling task.
4.3 Zero-Shot and Few-Shot Performance
@@
In Tab. 3, we present the zero-shot performance of various methods on LLaMA-2 7B, LLaMA-3 8B, and Qwen-2.5 7B. Our method consistently achieves the best average performance across all models. For LLaMA-2 7B, compared to weaker methods like ShortGPT and SliceGPT, our approach demonstrates significant advantages (ToMoE 50%: 60.72 vs. SliceGPT 40.84 and ShortGPT 70%: 47.07). The advantage against stronger baselines is also obvious. Although ModeGPT performs closer to ToMoE, the gap remains significant. With 60% active parameters, ToMoE is 3.14 times better than ModeGPT. For LLaMA-3 8B, the performance advantage of ToMoE is even larger, where it reduces 5% more active parameters than ModeGPT while still achieving a 2.71 performance gain. Furthermore, when removing 15% more active parameters compared to ShortGPT and SliceGPT, ToMoE exceeds their average performance by 18.93 and 16 points, respectively. For Qwen-2.5 7B, ToMoE significantly outperforms DISP-LLM, consistent with previous findings on other models. We further investigate the effect of the number of experts $N$ when 40% to 50% of the parameters are active. The results indicate that increasing the number of experts to 16 is beneficial. However, further increasing $N$ to 24 provides only marginal or no improvement, likely because a too-large number of experts burdens the learning process. Thus, we recommend choosing the number of experts $N$ to be smaller than 16.
In Tab. 4, our method demonstrated superb advantages compared to existing MoE construction methods. In “+fine-tuning" setting of LLaMA-MoE, the resulting model is trained for the same number of iterations as ToMoE for updating model weights. In Tab. 5 and Tab. 6, we further compare our method with CMoE, LLaMA-MoE, and LLaMA-MoE-v2, following the experimental settings in their papers. Our approach consistently surpasses all baselines while requiring significantly fewer tokens. Notably, in Tab. 6, when compared against the fully trained LLaMA-MoE-v2, our method achieves comparable performance even without additional fine-tuning of the model weights. In summary, our method shows that learning routers and experts together is a more promising solution compared to existing works.